# Multi-Modal and Multi-Agent Systems Meet Rationality: A Survey

### v2

@@ 4.3 Towards Grounding & Invariance & Independence from Irrelevant Contexts through Tool Utilization
Enabling agents to use tools also expands their bounded working memories, akin to retrieving external knowledge. Toolformer Schick et al. (2024) opens a new era that allows LLMs to use external tools via API calls following predefined syntax, effectively extending their capabilities beyond their intrinsic limitations and enforcing consistent and predictable outputs.
A multi-agent system can coordinate agents understanding when and which tool to use, which modality of information the tool should expect, how to call the corresponding API, and how to incorporate outputs from the API calls, which anchors subsequent reasoning processes with more accurate information beyond their parametric memory. For example, VisProg Gupta and Kembhavi (2023), ViperGPT Surís et al. (2023), and Parsel Zelikman et al. (2023) generate Python programs to reliably execute subroutines. Gupta and Kembhavi (2023); Surís et al. (2023) also invoke off-the-shelf models for multimodal assistance. Foundation models are not specifically trained for object detection or segmentation, so BuboGPT Zhao et al. (2023) and Multi-Agent VQA Jiang et al. (2024) call SAM Kirillov et al. (2023); Ren et al. (2024) as the tool. Jiang et al. (2024) finds $8.8\%$ of accuracy improvements compared to a single agent. Besides, BabyAGI Nakajima (2023), Chamelon Lu et al. (2024), AssistGPT Gao et al. (2023b), Avis Hu et al. (2024), ToolAlpaca Tang et al. (2023), MetaGPT Hong et al. (2023a), Agent LUMOS Yin et al. (2023), AutoAct Qiao et al. (2024), $\alpha$-UMi Shen et al. (2024a), and ConAgents Shi et al. (2024) harness compositional reasoning to enable generalized multi-agent systems with planning and modular tool-using capabilities in real-world scenarios.
Abstraction that Boils Down to Logical Essence
In most cases, tools require translating natural language queries into API calls with predefined syntax. Once the planning agent has determined the APIs and their input arguments, the original queries that may contain irrelevant contexts become invisible to the tools, and the tools will ignore any variance in the original queries as long as they share the equivalent underlying logic. Take Multi-Agent VQA Jiang et al. (2024) as an example. It has an LLM which provides only relevant object names rather than the whole visual question to the Grounded SAM Ren et al. (2024) component of the system acting as an object-detector. Similarly, the image editing tools in VisProg Gupta and Kembhavi (2023) only receive a fixed set of arguments translated from user queries to perform deterministic code executions. SeeAct Zheng et al. (2024a) as a Web agent explores vision-language models, ranking models, and a bounding box annotation tool to improve Web elements grounding from lengthy and noisy HTML codes. Consequently, using tools in a multi-agent system enhances the invariance and independence from irrelevant contexts, ensuring that their operations are streamlined and focused solely on necessary information.
4.4 Towards Orderability of Preferences & Invariance & Independence from Irrelevant Context through Neuro-Symbolic Reasoning
Alignment with Dual-Process Theories
Neural-symbolic reasoning is another promising approach to achieving consistent ordering of preferences and invariance by combining the strengths of languages and symbolic logic in a multi-agent system. Nye et al. (2021) bridge neural-symbolic reasoning with the dual-process theories. It posits that LLMs are predominantly single systems operating in a mode akin to System 1. To address this, it introduces a multi-agent system that includes System 2-like agents to break down complex tasks into symbolic programs, filter candidate responses generated the previous LLM, and compare them with a minimal world model to check consistency. Such a slower interaction between LLMs and neuro-symbolic modules enhances the coherence and explainablity of the overall reasoning process.
Coherent Orderability of Preference
A multi-agent system incorporating symbolic modules can not only understand language queries but also solve them with a level of consistency, providing a faithful and transparent reasoning process based on well-defined rules that adhere to logical principles, which is unachievable by LLMs alone. Logic-LM Pan et al. (2023) combines problem formulating, symbolic reasoning, and result interpreting agents, where the symbolic reasoner empowers LLMs with deterministic symbolic solvers to perform inference, ensuring a correct answer is consistently chosen. Its multi-agent framework also encourages self-refinement that modifies logical formulation errors using error messages from the symbolic reasoner as the feedback. Besides, SymbolicToM Sclar et al. (2023b) and KRISP Marino et al. (2021) construct explicit symbolic graphs and answer questions by retrieving nodes in the graph. Binder Cheng et al. (2022), Parsel Zelikman et al. (2023), LEFT Hsu et al. (2024), and Fang et al. (2024) decompose tasks into planning, parsing, and execution, where the symbolic reasoning agents can help maintain a coherent order of preferences among symbolic options in the system outputs. By skipping the symbolic module, Parsel Zelikman et al. (2023) observes a substantial performance drop of $19.5\%$. LEFT Hsu et al. (2024) also outperforms end-to-end baselines without symbolic programs by $3.85\%$ on average across multiple experiments. In more explicit scenarios, logical modules can directly compare the order of options represented as variables—such as “left" or “right" in relational logic Hsu et al. (2024)—rather than relying on a single LLM to generate responses indeterministically within the natural language space.
Abstraction that Boils Down to Logical Essence
Neuro-Symbolic modules typically expect standardized input formats Zelikman et al. (2023); Pan et al. (2023); Sclar et al. (2023b); Hsu et al. (2024); Fang et al. (2024); Yang et al. (2024); Subramanian et al. (2024), analogous to making API calls with external tools. This layer of abstraction enhances the independence from irrelevant contexts and maintains the invariance of LLMs when handling natural language queries. The only relevant factor is the parsed inputs into the predetermined neuro-symbolic programs. For instance, Ada Wong et al. (2023) introduces symbolic operators to abstract actions, ensuring that lower-level planning models are not compromised by irrelevant information in the queries and observations. Without the symbolic action library, a single LLM would frequently fail at grounding objects or obeying environmental conditions, resulting in a significant accuracy gap of approximately $59.0$-$89.0\%$.
4.5 Towards Orderability of Preferences & Invariance through Reflection, Debate, and Memory
Alignment with Dual-Process Theories
Due to the probabilistic outputs of LLMs, which resemble the rapid, non-iterative nature of human System 1 cognition, ensuring preference orderability and invariance is challenging. In contrast, algorithms that enable self-reflection and multi-agent systems that promote debate and consensus can help align outputs more closely with the deliberate and logical decision-making typical of System 2 processes, thus enhancing rational reasoning in agents.
Memory is one of the most fundamental cognitive processes that lead to reasoning, creativity, learning, and even self-consciousness in humans Solso and Kagan (1979); Craik and Lockhart (1972); Leydesdorff and Hodgkin (2017); Johnson-Laird (1983); Laird (2019); Sun (2001). In other words, even an artificial general intelligence (AGI) that lacks memory and forgets every past conversation after each query would find it hard to develop a coherent and rational decision-making.
Deliberation that Slows Down the “Thinking" Process
A narrow definition of the agent memory includes historical information within the same conversation Zhang et al. (2024b), like the multi-round self-reflection prompting strategies that encourage agents to critically evaluate their previous responses Shinn et al. (2024); Madaan et al. (2024); Wang et al. (2022b); Zhong et al. (2024); Lu et al. (2023). Historical conservations, usually from different or opposite perspectives that corrects initial errors, help produce a more error-prone and consistent final decision. Yao et al. (2024, 2023b); Besta et al. (2024) even support backtracking to further slow down the “thinking" process.
A broader definition of agent memory expands to historical information across multi-tasks and multi-agents Zhang et al. (2024b). Corex Sun et al. (2023a) finds that orchestrating multiple agents to work together yields better complex reasoning results, exceeding strong single-agent baselines Wang et al. (2022b) by an average of $1.1$-$10.6\%$. Retroformer Yao et al. (2023a) equips the single-agent Reflexion Shinn et al. (2024) algorithm with an additional LLM to generate verbal reinforcement cues and assist its self-improvement, enhancing accuracy by $1.0$-$20.9\%$. ChatEval Chan et al. (2023) introduces a multi-agent debate framework to mimic human annotators collaborating in robust answer evaluations. Its multi-agent approach achieves greater alignment with human preferences compared to single-agent evaluations, enhancing accuracy by $6.2\%$ for GPT-3.5 and $2.5\%$ for GPT-4, and an increase of $16.3\%$ and $10.0\%$ in average Spearman and Kendall-Tau correlations Zhong et al. (2022) with human judgements in GPT-4. MetaAgents Li et al. (2023j) effectively coordinates agents within task-oriented social contexts to achieve consistent behavior patterns, and the implementation of agent reflection in this system leads to a $21.0\%$ improvement in success rates.
@@ 4.5 Towards Orderability of Preferences & Invariance through Reflection, Debate, and Memory
LM vs LM Cohen et al. (2023), FORD Xiong et al. (2023), Multi-Agent Debate Liang et al. (2023); Du et al. (2023), DyLAN Liu et al. (2023d), and Khan et al. (2024) highlight the profound impact of multi-agent collaboration through cross-examination and debates. These studies demonstrate substantial improvements in performance when multiple agents are orchestrated to work in collaboration. Specifically, LM vs LM Cohen et al. (2023) illustrates how its multi-agent framework improves F1 scores by an average of $15.7\%$ compared to the single-agent baseline Yoshikawa and Okazaki (2023). FORD Xiong et al. (2023) reports an accuracy increase up to $4.9\%$ compared to a single LLM. Liang et al. (2023) indicates significant improvements in accuracy — $17.0\%$ for translation tasks and $16.0\%$ for reasoning tasks — by employing a multi-agent strategy, effectively bridging the performance gap between GPT-3.5 and GPT-4 by harnessing multi-agents. Du et al. (2023) finds that multi-agent debates not only enhance reasoning performance by $8.0$-$14.8\%$, but more importantly, increase factual accuracy by $7.2$-$15.9\%$. DyLAN Liu et al. (2023d) observes $3.5$-$4.1\%$ in accuracy improvements over single-agent execution. Multi-agent debating in Khan et al. (2024) also leads to more truthful answers, boosting single-agent baselines by $28.0\%$. Multi-Agent Collaboration Talebirad and Nadiri (2023), ChatDev Qian et al. (2023), AgentCF Zhang et al. (2023), AutoGen Wu et al. (2023), Social Learning Mohtashami et al. (2023), S3 Gao et al. (2023a), Ke et al. (2024), and Chern et al. (2024) continue to push the frontier of a multi-agent system’s applications beyond daily conversation to a versatile set of real-world task completions.
LLMs are also sensitive to prompt perturbations due to token bias and noises Sclar et al. (2023a). One of the most worrying examples are adversarial attacks Gehman et al. (2020); Ganguli et al. (2022); Du et al. (2022); Wei et al. (2024); Perez et al. (2022); Zou et al. (2023) through malicious prompt engineering, also named the Red Team Task. Chern et al. (2024) introduce a multi-agent debating approach with agents having harmless, neutral, or harmful agent intentions. It finds that multi-round multi-agent debate surpasses the self-reflection of a single agent, thus improving the invariance.
These collaborative approaches, in summary, allow each agent in a system to compare and rank its preference on choices from its own or from other agents through critical judgments. It helps enable the system to discern and output the most dominant decision as a consensus, thereby improving the orderability of preference. At the same time, through such a slow and critical thinking process, errors in initial responses or input prompts are more likely to be detected and corrected. Accumulated experience from past error planning contributes to a self-evolving process within the multi-agent system Zhang et al. (2024b), resulting in a final response or a consensus that is less sensitive to specific wording or token bias, moving the response towards better consistency and invariance.
@@ 5 Evaluating Rationality of Agents
The amount of studies for testing rationality in multi-modal and multi-agent systems remains scant, despite the growing interest in the field. While there are numerous reasoning benchmarks available, including commonsense reasoning (Talmor et al., 2019), logical reasoning Liu et al. (2021, 2023a), multi-hop reasoning (Yang et al., 2018), mathematical reasoning (Hendrycks et al., 2021), structured data reasoning Chen et al. (2020), conceptual reasoning Zhou et al. (2024), and general capabilities through multi-agent evaluations Rasheed et al. (2024); Ma et al. (2024); Wang et al. (2024b); Abdelnabi et al. (2023).
However, they do not directly measure rationality. Many of these benchmarks fail to prove whether reasoning is actually used in solving the tasks, leaving no guarantee that these tasks will be solved consistently when generalized to other representations or domains. Issues such as data contamination Magar and Schwartz (2022); Dong et al. (2024); Sainz et al. (2023); Jacovi et al. (2023) further compound the problem, as some benchmarks may inadvertently include the training data of these LLMs, leading to inflated performance scores. Hence, even though solid reasoning will imply rationality, existing approaches fall short in making the logic click. In this section, we point to several existing ingredients that can constitute the bread-and-butter of future generations of evaluation approaches for rationality.
@@
Testing Grounding against Hallucination
Information grounding is usually evaluated by the level of hallucination Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023), which can impact the rationality of agent systems. Multiple evaluation benchmarks targeting language-only dialogue have been proposed, such as BEGIN Dziri et al. (2022b), HaluEval Li et al. (2023f), DialFact Gupta et al. (2021), FaithDial Dziri et al. (2022a), AIS Rashkin et al. (2023), and others Zheng et al. (2023b); Das et al. (2023); Cao et al. (2021). In contrast, benchmarks on multi-agent frameworks beyond language dialogue or those involving multi-modalities are very limited. Liu et al. (2024a) moves beyond conversation to code generation; EureQA Li et al. (2023a) focuses on reasoning chains; and TofuEval Tang et al. (2024) evaluates hallucination in multi-domain summarization. Object hallucination Rohrbach et al. (2018); Biten et al. (2022), POPE Li et al. (2023h), and LLaVA-RLHF Sun et al. (2023b) are the few examples evaluating multi-modal hallucination. The community needs more hallucination benchmarks to quantitatively evaluate the extent to which multi-modal and multi-agents reduce hallucinations in comparison with baselines.
@@
There are almost no benchmarks for evaluating whether LLMs or agents have a consistent preference in the selection of available options. The Multiple Choice Problem (MCP) serves as a common testing ground. Zheng et al. (2023a) shows that LLMs are susceptible to changes in the positioning of options. Since the underlying logic remains the same, it also makes LLMs fail to pass the property of invariance. There are tons of MCP benchmarks PaperswithcodeMCQA , but solely on the accuracy of selections, overlooking the consistency of preference. However, Robinson et al. (2023) highlights that the Proportion of Plurality Agreement (PPA) offers a measure of order invariance that does not depend on the model’s ability to perform a task, suggesting a promising direction. We also need evaluations into vision and other modalities.
@@
The realization of language can be independent of its meaning Frege (1892); Wittgenstein (1953); Heineman (2023). Recent studies within the realm of data contamination investigate whether LLMs can generate consistent responses across different, yet inherently equivalent, framings of the same task. These studies introduce perturbations to the original task descriptions to assess whether LLMs’ responses will change significantly. Perturbation techniques include modifying instruction templates Weber et al. (2023), paraphrasing task descriptions (Yang et al., 2023; Ohmer et al., 2024), or altering the order of in-context learning exemplars Lu et al. (2021); Pecher et al. (2024). Specifically, approaches involve some versions of paraphrasing or permutation, such as changing the instruction templates Weber et al. (2023), rewording task descriptions (Yang et al., 2023; Ohmer et al., 2024; Wang et al., 2024b), translating the prompts into a different language (Ohmer et al., 2023, 2024; Xu et al., 2024a) and then back to the original language Yang et al. (2023), and making subtle changes to entities in task descriptions without affecting the logical structure, like altering names of the characters, numerical values in math problems, or locations of the events Wang et al. (2024b). Permutation also includes reordering in-context learning examples Lu et al. (2021); Pecher et al. (2024) and, in the case of multiple-choice questions, rearranging the options (Zong et al., 2023; Zheng et al., 2023a).
It is crucial to recognize that these perturbations are superficial; the altered task descriptions remain syntactically and semantically equivalent to their originals, although linguistic expressions or narratives may vary substantially. This observation highlights the need for developing methods that go beyond surface-level perturbations to effectively evaluate the robustness and invaraince of LLMs across diverse problem framings and modalities.
@@
We need more expansive perturbations to evaluate the independence from irrelevant context. Studies such as Shi et al. (2023), Wu et al. (2024), Liu et al. (2024d), and Yoran et al. (2023) have explored the phenomenon of “lost-in-context" by introducing random or misleading sentences into original problem statements. While earlier benchmarks like those from Weston et al. (2015), Sinha et al. (2019), Clark et al. (2020), and Webson and Pavlick (2021) have included irrelevant content, they have been predominantly limited to language modalities and single-agent systems. Recent benchmarks such as MileBench Song et al. (2024), Mementos Wang et al. (2024c), Seed-bench-2 Li et al. (2023b), and DEMON Li et al. (2023d) begin to evaluate multi-modal agents in long context or image sequences, where accurately responding to a specific question requires isolating only the relevant information from the long context window.

### v3

@@ 4.1 Towards Grounding & Invariance
Enabling agents to use tools also expands their bounded working memories, akin to retrieving external knowledge. Toolformer Schick et al. (2024) opens a new era that allows LLMs to use external tools via API calls following predefined syntax, effectively extending their capabilities beyond their intrinsic limitations and enforcing consistent and predictable outputs. A multi-agent system can coordinate agents understanding when and which tool to use, which modality of information the tool should expect, how to call the corresponding API, and how to incorporate outputs from the API calls, which anchors subsequent reasoning processes with more accurate information beyond their parametric memory. For example, VisProg Gupta and Kembhavi (2023), ViperGPT Surís et al. (2023), and Parsel Zelikman et al. (2023) generate Python programs to reliably execute subroutines. Gupta and Kembhavi (2023); Surís et al. (2023) also invoke off-the-shelf models for multimodal assistance. For more examples, see Appendix C.
4.2 Towards Rationality through Deliberation
Memory is one of the most fundamental cognitive processes that lead to reasoning, creativity, learning, and even self-consciousness in humans Solso and Kagan (1979); Craik and Lockhart (1972); Leydesdorff and Hodgkin (2017); Johnson-Laird (1983); Laird (2019); Sun (2001). Any system that lacks the ability to retain information from previous interactions would struggle to make coherent and rational decisions in the long run. In a narrow sense, agent memory includes historical information within the same conversation Zhang et al. (2024b). This allows for deliberation, which is the slower, iterative reasoning process to carefully consider information and options in order to arrive at more rational and well-reasoned decisions.
Although deliberation may increase the likelihood of reaching more rational decisions, there is no inherent guarantee for rationality via this approach. The quality of the decision ultimately depends on the accuracy and relevance of the grounded information, as well as the soundness of the reasoning process. Biases, incomplete information, and flawed logic can still lead to irrational conclusions even with deliberation. Nonetheless, multiple works have shown that the increase in likelihood of rational decisions through deliberation is significant and beneficial. For example, multi-round self-reflection prompting strategies that encourage agents to critically evaluate their previous responses Shinn et al. (2024); Madaan et al. (2024); Wang et al. (2022b); Zhong et al. (2024); Lu et al. (2023).
In a broader context, for multi-agent systems, agent memory expands to include historical information across multiple tasks and agents Zhang et al. (2024b). This shared memory enables collective deliberation among agents via collaboration, cross-examination, and debates. By leveraging the collective knowledge and experiences of multiple agents, the system can arrive at more comprehensive and robust solutions to complex problems.
@@ 4.2 Towards Rationality through Deliberation
LM vs LM Cohen et al. (2023), FORD Xiong et al. (2023), Multi-Agent Debate Liang et al. (2023); Du et al. (2023), DyLAN Liu et al. (2023c), and Khan et al. (2024) highlight the profound impact of multi-agent collaboration through cross-examination and debates. Specifically, LM vs LM Cohen et al. (2023) illustrates how its multi-agent framework improves F1 scores by an average of $15.7\%$ compared to the single-agent baseline Yoshikawa and Okazaki (2023). FORD Xiong et al. (2023) reports an accuracy increase up to $4.9\%$ compared to a single LLM. Liang et al. (2023) indicates significant improvements in accuracy — $17.0\%$ for translation tasks and $16.0\%$ for reasoning tasks — by employing a multi-agent strategy, effectively bridging the performance gap between GPT-3.5 and GPT-4 by harnessing multi-agents. Du et al. (2023) finds that multi-agent debates not only enhance reasoning performance by $8.0$-$14.8\%$, but more importantly, increase factual accuracy by $7.2$-$15.9\%$. DyLAN Liu et al. (2023c) observes $3.5$-$4.1\%$ in accuracy improvements over single-agent execution. All these approaches enhance the system’s capability to capture initial errors, improve factuality in reasoning, and achieve final consensus with fewer inconsistencies. We discuss more examples in Appendix D.1. We also talk about collaboration against jailbreaking in Appendix D.2 and multi-agent evaluation methods in Appendix D.3.
4.3 Towards Rationality through Abstraction
Independence from irrelevant contexts, invariance, and orderability of preferences can be achieved simultaneously through the use of tools and neuro-symbolic reasoning, because these approaches translate natural language queries into standardized formats like API calls or symbolic representations, which abstract away extraneous details, focus only on the underlying logic necessary for the task at hand, and enable consistent and deterministic processing of the input.
Independence from Irrelevant Contexts
In most cases, tools require translating natural language queries into API calls with predefined syntax. Once the APIs and their input arguments are determined, the tools will ignore any irrelevant context in the original queries, as long as the queries share the same underlying logic necessary for the inputs. Take Multi-Agent VQA Jiang et al. (2024b) as an example. In this system, a language model provides only the relevant object names to the Grounded SAM Ren et al. (2024) component, which functions as an object detector, rather than passing the entire visual question. Other similar examples are discussed in Appendix C.
Neuro-symbolic reasoning is an approach that combines neural networks with symbolic systems, such as explicit knowledge representation and logical reasoning. A multi-agent system incorporating symbolic modules can not only understand language queries but also solve them with a level of consistency, providing a faithful and transparent reasoning process based on well-defined rules that adhere to logical principles, which is unachievable by a single language model.
Analogous to the external tool utilization, neuro-symbolic modules in a multi-agent system expect standardized input formats Zelikman et al. (2023); Pan et al. (2023); Sclar et al. (2023b); Hsu et al. (2024); Fang et al. (2024); Yang et al. (2024); Subramanian et al. (2024). The only relevant factor in this process is the parsed inputs into the predetermined neuro-symbolic programs. For instance, Ada Wong et al. (2023) introduces symbolic operators to abstract actions, ensuring that lower-level planning models are not compromised by irrelevant information in the queries and observations. Without the symbolic action library, a single LLM would frequently fail at grounding objects or obeying environmental conditions, resulting in a significant accuracy gap of approximately $59.0$-$89.0\%$.
Invariance
The abstraction provided by symbolic representations also allows the multi-agent system to solve language queries with invariance. For example, Logic-LM Pan et al. (2023) combines problem formulating, symbolic reasoning, and result interpreting agents, where the symbolic reasoner empowers LLMs with deterministic symbolic solvers to perform inference, ensuring a correct answer is consistently chosen. Its multi-agent framework also encourages self-refinement that modifies logical formulation errors using error messages from the symbolic reasoner as the feedback. For more examples, see Appendix E.
Orderability of Preferences
In explicit scenarios, logical modules can directly compare the order of options represented as variables—such as “left" or “right" in relational logic Hsu et al. (2024)—rather than relying on a single LLM to generate responses indeterministically within the natural language space. In more complex situations, systems like Binder Cheng et al. (2022), Parsel Zelikman et al. (2023), LEFT Hsu et al. (2024), and Fang et al. (2024) decompose tasks into planning, parsing, and execution, where the symbolic reasoning agents can help maintain a coherent order of preferences among symbolic options in the system outputs. By skipping the symbolic module, Parsel Zelikman et al. (2023) observes a substantial performance drop of $19.5\%$. LEFT Hsu et al. (2024) also outperforms end-to-end baselines without symbolic programs by $3.85\%$ on average across multiple experiments.
Recent work has also explored applying expected utility theory Von Neumann and Morgenstern (2007) to improve the decision-making capabilities of language models. For example, DeLLMa Liu et al. (2024e) decomposes complex decision problems into subtasks, assigns utilities to different outcomes, and selects actions that maximize expected utility.
@@ 5 Evaluating Rationality of Agents
The amount of studies for testing rationality in multi-modal and multi-agent systems remains scant, despite the growing interest in the field. While there are numerous reasoning benchmarks available (Talmor et al., 2019; Liu et al., 2021; Yang et al., 2018; Hendrycks et al., 2021), they do not directly measure rationality. Many of these benchmarks fail to prove whether reasoning is actually used in solving the tasks, leaving no guarantee that these tasks will be solved consistently when generalized to other representations or domains. Issues such as data contamination Magar and Schwartz (2022); Dong et al. (2024); Sainz et al. (2023); Jacovi et al. (2023) further compound the problem, as some benchmarks may inadvertently include the training data of these LLMs, leading to inflated performance scores. Hence, even though solid reasoning will imply rationality, existing approaches fall short in making the logic click. In this section, we point to several existing ingredients that can constitute the bread-and-butter of future generations of evaluation approaches for rationality.
@@
Testing against Hallucination
Information grounding is usually evaluated by the level of hallucination Bang et al. (2023); Guerreiro et al. (2023); Huang et al. (2023). Multiple evaluation benchmarks targeting language-only dialogue have been proposed, such as BEGIN Dziri et al. (2022b), HaluEval Li et al. (2023e), DialFact Gupta et al. (2021), FaithDial Dziri et al. (2022a), AIS Rashkin et al. (2023), and others Zheng et al. (2023b); Das et al. (2023); Cao et al. (2021). In contrast, benchmarks on multi-agent frameworks or those involving multi-modalities beyond language dialogue are very limited. We find that Liu et al. (2024a) moves beyond conversation to code generation, EureQA Li et al. (2023a) focuses on reasoning chains, and TofuEval Tang et al. (2024) evaluates hallucination in multi-domain summarization. Object hallucination Rohrbach et al. (2018); Biten et al. (2022), POPE Li et al. (2023g), and LLaVA-RLHF Sun et al. (2023b) are the few examples evaluating multi-modal hallucination. The community needs more hallucination benchmarks to quantitatively evaluate the extent to which multi-modal and multi-agents reduce hallucinations in comparison with baselines.
@@
There are almost no benchmarks for evaluating whether LLMs or agents have a consistent preference in the selection of available options. The Multiple Choice Problem (MCP) serves as a common testing ground. Zheng et al. (2023a) shows that LLMs are susceptible to changes in the positioning of options. Since the underlying logic remains the same, it also makes LLMs fail to pass the property of invariance. Although there are many MCP benchmarks PaperswithcodeMCQA , they focus on the accuracy of selections and overlook the consistency of preference. However, Robinson et al. (2023) highlights that the Proportion of Plurality Agreement (PPA) offers a measure of order invariance that does not depend on the model’s ability to perform a task, suggesting a promising direction.
@@
Recent studies concerning data contamination investigate whether LLMs can generate consistent responses across different, yet inherently equivalent, framing of the same task. These studies introduce perturbations to the original task descriptions to assess whether LLMs’ responses will change significantly. Perturbation techniques include modifying instruction templates Weber et al. (2023), paraphrasing task descriptions (Yang et al., 2023; Ohmer et al., 2024), or altering the order of in-context learning exemplars Lu et al. (2021); Pecher et al. (2024). For more details on these techniques, we refer the reader to Appendix F.2. It is crucial to recognize that these perturbations are superficial: the altered task descriptions remain syntactically and semantically equivalent to their originals, although linguistic expressions or narratives may vary substantially. Methods that go beyond surface-level perturbations are needed to evaluate the robustness and invariance of LLMs across diverse problem framings and modalities effectively.
@@
Studies such as Shi et al. (2023), Wu et al. (2024), Liu et al. (2024d), and Yoran et al. (2023) have explored the phenomenon of “lost-in-context" by introducing random or misleading sentences into original problem statements. While earlier benchmarks like those from Weston et al. (2015), Sinha et al. (2019), Clark et al. (2020), and Webson and Pavlick (2021) have included irrelevant content, they have been predominantly limited to language modalities and single-agent systems. Recent benchmarks such as MileBench Song et al. (2024), Mementos Wang et al. (2024c), Seed-bench-2 Li et al. (2023b), and DEMON Li et al. (2023c) begin to evaluate multi-modal agents in long context or image sequences, where accurately responding to a specific question requires isolating only the relevant information from the long context window.